# CUQDS: Conformal Uncertainty Quantification under Distribution Shift for Trajectory Prediction

### v2

@@ Abstract
Trajectory prediction models that can infer both future trajectories and their associated uncertainties of the target vehicles is crucial for safe and robust navigation and path planning of autonomous vehicles. However, the majority of existing trajectory prediction models have neither considered reducing the uncertainty as one objective during the training stage nor provided reliable uncertainty quantification during inference stage, especially under potential distribution shift. Therefore, in this paper, we propose the Conformal Uncertainty Quantification under Distribution Shift framework, CUQDS, to quantify the uncertainty of the predicted trajectories of existing trajectory prediction models under potential data distribution shift, while improving the prediction accuracy of the models and reducing the estimated uncertainty during the training stage. Specifically, CUQDS includes 1) a learning-based Gaussian process regression module that models the output distribution of the base model (any existing trajectory prediction neural networks) and reduces the estimated uncertainty by an additional loss term, and 2) a statistical-based Conformal P control module to calibrate the estimated uncertainty from the Gaussian process regression module in an online setting under potential distribution shift between training and testing data. Experimental results on various state-of-the-art methods using benchmark motion forecasting datasets demonstrate the effectiveness of our proposed design.
1 Introduction
@@
Accurately and efficiently predicting the future trajectories of the target vehicles is pivotal for ensuring the safety of path planning and navigation of autonomous systems in dynamic environments Hu et al. (2023); Kedia et al. (2023). However, trajectory uncertainly due to the changing external surrounding environments (e.g., road networks, neighborhood vehicles, passengers, and so on) or intrinsic intention changes of drivers, can lead to both trajectory distribution shift that change over time and overconfidence in the output of trajectory prediction models. It remains challenging to predict the distribution of the future trajectories of the target vehicles (rather than focusing solely on point estimates), and to quantify the uncertainty of the output trajectories under potential distribution shift between training and testing data.
@@
Various uncertainty quantification methods have been proposed for trajectory prediction. One popular approach is to estimate the distribution (e.g., Laplace distribution or Gaussian distribution) of future trajectories of target vehicles by direct modeling methods Zhou et al. (2022); Mao et al. (2023); Zhu et al. (2023); Salzmann et al. (2020); Chen et al. (2023); Tang et al. (2021). However, this method often overlooks the impact of model limitation (the restricted ability of models to represent the real-time trajectories data) and the consequence overconfident uncertainty estimation in the inference stage. Another popular approach is to calibrate the preliminary estimated uncertainty in the inference stage by the statistical-based methods. Specifically, methods such as split conformal prediction (CP) Shafer and Vovk (2008); Lindemann et al. (2023); Lekeufack et al. (2023) provide confidence intervals that guarantee to contain the ground truth target data with a predefined probability. However, they are not applicable to the situations when there is distribution shift between the training and testing data.
@@
In this paper, we propose CUQDS, i.e., Conformal Uncertainty Quantification under Distribution Shift framework, to quantify the output uncertainty of existing trajectory prediction models under distribution shift, while improving the prediction accuracy of the models and reducing the estimated uncertainty during the training stage. The proposed CUQDS framework integrates a learning-based Gaussian process regression Rasmussen et al. (2006) module with a statistics-based conformal P control module Angelopoulos et al. (2023). Specifically, in the training stage, CUQDS adopts the Gaussian process regression module to estimate the output distribution of base model. The variance of the output distribution quantifies the uncertainty of the predicted trajectories. We further use the output distribution to construct uncertainty interval that guarantees to cover the true target trajectory with a predefined probability in the long-run. As shown in Fig. 1, a good uncertainty interval should tend to be narrow when the predicted trajectory closely aligns with the ground truth trajectory (e.g., vehicle A), whereas it should be widen when the output trajectory deviates (e.g., vehicle B). To achieve this, we introduce additional loss term to reduce the estimated uncertainty in the training stage, ensuring narrow uncertainty interval while covering the true target trajectory with a predefined probability. In the testing or inference stage, CUQDS calibrates the estimated output uncertainty from the Gaussian process regression module upon potential distribution shift by a statistics-based conformal P control module. Different from standard conformal prediction method and its variants who use fixed conformal quantile in the inference stage and violate the distribution shift assumption, we first initialize the conformal quantile using the validation data and update it after every prediction step during the inference stage.
@@
    We validate the effectiveness of our proposed framework on the Argoverse 1 motion forecasting dataset Chang et al. (2019) and five state-of-the-art baselines. Compared to base models without our CUQDS, the experiment results show that our approach improves prediction accuracy by an average of 7.07% and reduce uncertainty of predicted trajectory by an average of 25.41% through incorporation with our CUQDS.
2 Related Work
@@
Modeling the distribution of the trajectory of the target vehicles, instead of solely focusing on the point estimation of the future trajectory, proves to be an efficient approach Chai et al. (2019); Deo and Trivedi (2018); Phan-Minh et al. (2020); Zeng et al. (2021) to avoid missing potential behavior in trajectory prediction methods. Early works sample multiple potential future trajectories Liang et al. (2020); Ngiam et al. (2021); Gupta et al. (2018); Rhinehart et al. (2018, 2019); Tang and Salakhutdinov (2019) to approximate the predicted trajectory distribution. However, these approaches still rely on limited point estimation and suffer from overconfidence upon trajectory prediction in real-world settings. To mitigate these issues, recent works directly estimate the distribution of future trajectories by fitting the (mixture) distribution models based on the embedded features of input trajectories Zhou et al. (2022); Varadarajan et al. (2022); Cui et al. (2019); Shi et al. (2023). However, these works mostly train the (mixture) distribution models by the MLP-based predictor in the last layer of the model. Such designs fail to estimate an accurate enough distribution of future trajectories under data distribution shift.
@@
Uncertainty quantification in trajectory prediction is challenging and usually solved by two categories of methods: direct modeling and statistical-based methods. Specifically, direct modeling assumes the target data follows specific distribution, designs the learning-based neural networks for distribution estimation, and introduces the corresponding loss function to model the uncertainty directly. To achieve this, existing works usually assume the data follows Gaussian distribution Mao et al. (2023) or Laplace distribution Gu et al. (2024) and model the corresponding distribution by learning-based designs. However, estimating rigorous uncertainty by direct modeling can be challenging, as the model may easily overfit the training dataset such that invalid in the inference stage under data distribution shift. Conformal prediction Shafer and Vovk (2008); Angelopoulos et al. (2023) stands out as a method in statistical-based inference that has proven effective in constructing predictive sets, ensuring a certain probability of covering the true target values Ivanovic et al. (2022); Xu et al. (2014). However, the assumption of identical distribution across all datasets invalidates the standard conformal prediction under distribution shift. Consequently, this necessitates modifications Xu and Xie (2023) in conformal prediction to enable timely calibration of model outputs when confronted with distribution shift.
@@
3 Methodology
3.1 Problem Formulation
Suppose we have a training dataset $\mathcal{D}_{1}=\{(x_{i},y_{i})\}_{i=1}^{N_{1}}=\{(x^{t},y^{t})\}_{t=L-1}^{T_{1}}$, a validation dataset $\mathcal{D}_{2}=\{(x_{i},y_{i})\}_{i=1}^{N_{2}}=\{(x^{t},y^{t})\}_{t=L-1}^{T_{2}}$, and a testing dataset $\mathcal{\tilde{D}}=\{(\tilde{x}_{i},\tilde{y}_{i})\}_{i=1}^{N_{3}}=\{(\tilde{x}^{t},\tilde{y}^{t})\}_{t=L-1}^{T_{3}}$. $N_{1}$, $N_{2}$, and $N_{3}$ are the number of data samples, and $T_{1}$, $T_{2}$, and $T_{3}$ are the time periods of each dataset. $T_{1}\bigcap T_{2}\bigcap T_{3}=\varnothing$. As shown above, we use two ways to represent each dataset, where $(x_{i},y_{i})$ (or $(\tilde{x}_{i},\tilde{y}_{i})$) denotes the random input-output trajectory pair within each dataset indexed by $i$, and $(x^{t},y^{t})$ (or $(\tilde{x}^{t},\tilde{y}^{t})$) is time series data and represents the input-output trajectory pair of the current time step $t$. Both $x_{i}$ and $x^{t}$ denote the input historical trajectory during the past $L$ historical time steps and are sampled from domain $\mathcal{X}\in\mathbb{R}^{L\times D}$, and $y_{i}$ and $y^{t}$ denote the corresponding target trajectory during the $J$ following time steps from domain $\mathcal{Y}\in\mathbb{R}^{J\times D}$. $D$ is the dimension of target features. We will omit the index $i$ or $t$ when there is no conflict.
We assume that we have a trajectory prediction model $f$ with parameters $\theta$. We call this model $f_{\theta}$ a base model and it can be implemented as different structures of deep neural network Zhou et al. (2022); Liu et al. (2021); Zhou et al. (2023); Liang et al. (2020); Zhong et al. (2022). The model $f_{\theta}$ is trained using the dataset $\mathcal{D}$, where the target trajectory, denoted as $Y$, is conditioned on input $x$ and the dataset $\mathcal{D}$. We model the conditional distribution of $Y$ given $x$ and $\mathcal{D}$ as a Gaussian distribution, where $Y|x,\mathcal{D}\sim\mathcal{N}\left(\mu(x),\sigma^{2}(x)\right)$. In this formulation, $\mu(x)$ and $\sigma^{2}(x)$ are functions that map the input $x$ to the mean and variance of the Gaussian distribution, respectively. We treat the predicted trajectory of the base model $f_{\theta}$ as the mean $\mu(x)$ directly. Our main task in this study is to estimate the variance $\sigma^{2}(x)$ to denote the output uncertainty.
@@ 3.1 Problem Formulation
The loss $\mathcal{L}$ is a weighted combination of base model’ loss $\mathcal{L}_{1}(\theta)$ and Gaussian processing regression model’s loss $\mathcal{L}_{2}(\omega,\theta)$.
@@
3.2 Gaussian Process Regression Module
@@
Existing literature of trajectory prediction models Zhou et al. (2022, 2023); Liang et al. (2020); Zhong et al. (2022) focuses on providing point estimates of future trajectories. However, the uncertainty of the future trajectories due to the changing environment or the intrinsic intention changes of drivers can lead to significant distribution shift and overconfident trajectory prediction. Such distribution shift and overconfident prediction can greatly impact the subsequent decision-making processes, such as robust path planning Hu et al. (2023); Kedia et al. (2023). Hence, we propose to consider output uncertainties from both the modeling limitation of the model $f_{\theta}$ upon current trajectory inputs and the noise inherent in the trajectory data.
1
@@
Input: a base model $f(\cdot)$ with an initialized parameter $\theta$ and a covariance function $k(\cdot)$ with an initialized parameter $\omega$, error rate $\alpha\in[0,1]$, score function $s(\cdot)$, total number of training epochs $epo$.
Data: training dataset $D_{1}=\{x_{i},y_{i}\}_{i=1}^{N_{1}}$, validation dataset $D_{2}=\{x^{t},y^{t}\}_{t=1}^{T_{2}}$.
@@
Output: well-trained base model $f_{\theta}(\cdot)$, updated statistic $\hat{q}^{1}$, score set $S={\{s\left(x^{t},y^{t}\right)\}}_{t=L-1}^{T_{2}}$ and error set $E={\{e^{t}\}}_{t=L-1}^{T_{2}}$ estimated from the validation data.

### v3

@@ Abstract
Trajectory prediction models that can infer both future trajectories and their associated uncertainties of the target vehicles are crucial for safe and robust navigation and path planning of autonomous vehicles. However, the majority of existing trajectory prediction models have neither considered reducing the uncertainty as one objective during the training stage nor provided reliable uncertainty quantification during inference stage, especially under potential distribution shift. Therefore, in this paper, we propose the Conformal Uncertainty Quantification under Distribution Shift framework, CUQDS, to quantify the uncertainty of the predicted trajectories of existing trajectory prediction models under potential data distribution shift, while improving the prediction accuracy of the models and reducing the estimated uncertainty during the training stage. Specifically, CUQDS includes 1) a learning-based Gaussian process regression module that models the output distribution of the base model (any existing trajectory prediction neural networks) and reduces the estimated uncertainty by an additional loss term, and 2) a statistical-based Conformal P control module to calibrate the estimated uncertainty from the Gaussian process regression module in an online setting under potential distribution shift between training and testing data. Experimental results on various state-of-the-art methods using benchmark motion forecasting datasets demonstrate the effectiveness of our proposed design.
Introduction
@@
Accurately and efficiently predicting the future trajectories of the target vehicles is pivotal for ensuring the safety of path planning and navigation of autonomous systems in dynamic environments (Hu et al. 2023; Kedia, Zhou, and Karumanchi 2023). However, trajectory uncertainly due to the changing external surrounding environments (e.g., road networks, neighborhood vehicles, passengers, and so on) or intrinsic intention changes of drivers, can lead to both trajectory distribution shift that change over time and overconfidence in the output of trajectory prediction models. It remains challenging to predict the distribution of the future trajectories of the target vehicles (rather than focusing solely on point estimates), and to quantify the uncertainty of the output trajectories under potential distribution shift between training and testing data.
@@
Various uncertainty quantification methods have been proposed for trajectory prediction. One popular approach is to estimate the distribution (e.g., Laplace distribution or Gaussian distribution) of future trajectories of target vehicles by direct modeling methods (Zhou et al. 2022; Mao et al. 2023; Zhu et al. 2023; Salzmann et al. 2020; Chen et al. 2023; Tang et al. 2021). However, this method often overlooks the impact of model limitation (the restricted ability of models to represent the real-time trajectories data) and the consequence overconfident uncertainty estimation in the inference stage. Another popular approach is to calibrate the preliminary estimated uncertainty in the inference stage by the statistical-based methods. Specifically, methods such as split conformal prediction (CP) (Shafer and Vovk 2008; Lindemann et al. 2023; Lekeufack et al. 2023) provide confidence intervals that guarantee to contain the ground truth target data with a predefined probability. However, they are not applicable to the situations when there is distribution shift between the training and testing data.
@@
In this paper, we propose CUQDS, i.e., Conformal Uncertainty Quantification under Distribution Shift framework, to quantify the output uncertainty of existing trajectory prediction models under distribution shift, while improving the prediction accuracy of the models and reducing the estimated uncertainty during the training stage. The proposed CUQDS framework integrates a learning-based Gaussian process regression (Rasmussen, Williams et al. 2006) module with a statistics-based conformal P control module (Angelopoulos, Candes, and Tibshirani 2023). Specifically, in the training stage, CUQDS adopts the Gaussian process regression module to estimate the output distribution of base model. The variance of the output distribution quantifies the uncertainty of the predicted trajectories. We further use the output distribution to construct uncertainty interval that guarantees to cover the true target trajectory with a predefined probability in the long-run. As shown in Fig. 1, a good uncertainty interval should tend to be narrow when the predicted trajectory closely aligns with the ground truth trajectory (e.g., vehicle A), whereas it should be widen when the output trajectory deviates (e.g., vehicle B). To achieve this, we introduce additional loss term to reduce the estimated uncertainty in the training stage, ensuring narrow uncertainty interval while covering the true target trajectory with a predefined probability. In the inference stage, CUQDS calibrates the estimated output uncertainty upon potential distribution shift by a statistics-based conformal P control module. Different from standard conformal prediction method and its variants who use fixed conformal quantile in the inference stage and violate the distribution shift assumption, we first initialize the conformal quantile using the validation data and update it after every prediction step during the inference stage.
@@
    We validate the effectiveness of our proposed framework on the Argoverse 1 motion forecasting dataset (Chang et al. 2019) and five state-of-the-art baselines. Compared to base models without our CUQDS, the experiment results show that our approach improves prediction accuracy by an average of 7.07% and reduce uncertainty of predicted trajectory by an average of 25.41%.
Related Work
@@
Modeling the distribution of the trajectory of the target vehicles, instead of solely focusing on the point estimation of the future trajectory, proves to be an efficient approach (Chai et al. 2019; Deo and Trivedi 2018; Phan-Minh et al. 2020; Zeng et al. 2021) to avoid missing potential behavior in trajectory prediction methods. Early works sample multiple potential future trajectories (Liang et al. 2020; Ngiam et al. 2021; Gupta et al. 2018; Rhinehart, Kitani, and Vernaza 2018; Rhinehart et al. 2019; Tang and Salakhutdinov 2019) to approximate the predicted trajectory distribution. However, these approaches still rely on limited point estimation and suffer from overconfidence upon trajectory prediction in real-world settings. To mitigate these issues, recent works directly estimate the distribution of future trajectories by fitting the (mixture) distribution models based on the embedded features of input trajectories (Zhou et al. 2022; Varadarajan et al. 2022; Cui et al. 2019; Shi et al. 2023). However, these works mostly train the (mixture) distribution models by the MLP-based predictor in the last layer of the model. Such designs fail to estimate an accurate enough distribution of future trajectories under data distribution shift.
@@
Uncertainty quantification in trajectory prediction is challenging and usually solved by two categories of methods: direct modeling and statistical-based methods. Specifically, direct modeling assumes the target data follows specific distribution, designs the learning-based neural networks for distribution estimation, and introduces the corresponding loss function to model the uncertainty directly. To achieve this, existing works usually assume the data follows Gaussian distribution (Mao et al. 2023) or Laplace distribution (Gu et al. 2024) and model the corresponding distribution by learning-based designs. However, estimating rigorous uncertainty by direct modeling can be challenging, as the model may easily overfit the training dataset such that invalid in the inference stage under data distribution shift. Conformal prediction (Shafer and Vovk 2008; Angelopoulos, Candes, and Tibshirani 2023) stands out as a method in statistical-based inference that has proven effective in constructing predictive sets, ensuring a certain probability of covering the true target values (Ivanovic et al. 2022; Xu et al. 2014; Sun et al. 2024; Xu et al. 2023; Stankeviciute, M Alaa, and van der Schaar 2021). However, the assumption of identical distribution across all datasets invalidates the standard conformal prediction under distribution shift. To address this, adaptive conformal prediction methods (Gibbs and Candes 2021; Stanton, Maddox, and Wilson 2023) have been proposed to achieve the desired coverage frequency under distribution shift by leveraging the distance between new queries and the training dataset. However, these approaches either lack base model uncertainty estimation or fail to incorporate mechanisms for reducing uncertainty during the training process, as achieved by our proposed CUQDS.
@@
Methodology
Problem Formulation
Suppose we have a training dataset $\mathcal{D}_{1}=\{(x_{i},y_{i})\}_{i=1}^{N_{1}}=\{(x^{t},y^{t})\}_{t=L-1}^{T_{1}}$, a validation dataset $\mathcal{D}_{2}=\{(x_{i},y_{i})\}_{i=1}^{N_{2}}=\{(x^{t},y^{t})\}_{t=L-1}^{T_{2}}$, and a testing dataset $\mathcal{\tilde{D}}=\{(\tilde{x}_{i},\tilde{y}_{i})\}_{i=1}^{N_{3}}=\{(\tilde{x}^{t},\tilde{y}^{t})\}_{t=L-1}^{T_{3}}$. $N_{1}$, $N_{2}$, and $N_{3}$ are the number of data samples, and $T_{1}$, $T_{2}$, and $T_{3}$ are the time periods of each dataset. $T_{1}\bigcap T_{2}\bigcap T_{3}=\varnothing$. In this paper, we assume that training/validation datasets are drawn from the same distribution, while the distribution of testing dataset shifts over time. As shown above, we use two ways to represent each dataset, where $(x_{i},y_{i})$ (or $(\tilde{x}_{i},\tilde{y}_{i})$) denotes the random input-output trajectory pair within each dataset indexed by $i$, and $(x^{t},y^{t})$ (or $(\tilde{x}^{t},\tilde{y}^{t})$) is time series data and represents the input-output trajectory pair of the current time step $t$. Both $x_{i}$ and $x^{t}$ denote the input historical trajectory during the past $L$ historical time steps and are sampled from domain $\mathcal{X}\in\mathbb{R}^{L\times D}$, and $y_{i}$ and $y^{t}$ denote the corresponding target trajectory during the $J$ following time steps from domain $\mathcal{Y}\in\mathbb{R}^{J\times D}$. $D$ is the dimension of target features. We will omit the index $i$ or $t$ when there is no conflict.
We assume that we have a trajectory prediction model $f$ with parameters $\theta$. We call this model $f_{\theta}$ a base model and it can be implemented as different structures of deep neural network (Zhou et al. 2022; Liu et al. 2021; Zhou et al. 2023; Liang et al. 2020; Zhong et al. 2022). The model $f_{\theta}$ is trained using the dataset $\mathcal{D}$, where the target trajectory, denoted as $Y$, is conditioned on input $x$ and the dataset $\mathcal{D}$. We model the conditional distribution of $Y$ given $x$ and $\mathcal{D}$ as a Gaussian distribution, where $Y|x,\mathcal{D}\sim\mathcal{N}\left(\mu(x),\sigma^{2}(x)\right)$. In this formulation, $\mu(x)$ and $\sigma^{2}(x)$ are functions that map the input $x$ to the mean and variance of the Gaussian distribution, respectively. We treat the predicted trajectory of the base model $f_{\theta}$ as the mean $\mu(x)$ directly. Our main task in this study is to estimate the variance $\sigma^{2}(x)$ to denote the output uncertainty.
@@
The loss $\mathcal{L}$ is a weighted combination of base model’s loss $\mathcal{L}_{1}(\theta)$ and Gaussian processing regression model’s loss $\mathcal{L}_{2}(\omega,\theta)$.
@@
Preliminary
$\bullet$ Gaussian Process Regression often acts as the surrogate model to estimate the output distribution of existing models (Erlygin et al. 2023). It introduces additional loss term to guild the learning process of existing models and reduce the output uncertainty. Gaussian process regression assumes any combinations of data samples follows different join distribution and thus good at capturing the nonlinear relationship among time series data samples.
Given an existing deep neural network model $f_{\theta}$, Gaussian process regression views this model as a black box and aims to estimate the output distribution $(Y|x,\theta\sim\mathcal{N}\left(\hat{\mu}\left(x\right),\hat{{\sigma}}^{2}\left(x\right)\right))$ corresponding to the inputs $x$. We treat the output of model $f_{\theta}\left(\cdot\right)$ on $x$ as the mean value $\hat{\mu}\left(x\right)$ of the estimated output distribution and the correspond variance $\hat{{\sigma}}^{2}\left(x\right)$ as the uncertainty of $f_{\theta}\left(x\right)$. Gaussian process regression also assumes that the observed values $y$ differ from the function values $f_{\theta}(x)$ by additive noise $\epsilon$, and further assumes that this noise follows an independent, identically distributed Gaussian distribution with zero mean and variance of ${\sigma}_{\epsilon}^{2}$. Now the prediction process is formulated as $Y|_{x}=\hat{\mu}(x)+\epsilon$, where $(x,y)$ is any data samples from $\mathcal{D}$. Then we have:
where $k(\cdot,\cdot)$ is the user-defined kernel function or the covariance function. For example, $k\left(x,x^{\prime}\right)$ quantifies the similarity between trajectory input data $x$ and $x^{\prime}$.
$\bullet$ Conformal Prediction (Shafer and Vovk 2008; Angelopoulos, Candes, and Tibshirani 2023) is a statistical-based method to construct predictive sets for any model, ensuring a certain probability of covering the true target values. It assumes identical distribution across all datasets. Given the base model $f_{\theta}$ whose output distribution follows $(Y|x,\theta\sim\mathcal{N}\left(\hat{\mu}\left(x\right),\hat{{\sigma}}^{2}\left(x\right)\right))$, conformal prediction converts the heuristic notion of uncertainty (e.g. the estimated standard deviation $\hat{{\sigma}}\left(x\right)$) into rigorous prediction intervals of the form $\left(\hat{\mu}\left(x\right)\pm q\hat{{\sigma}}\left(x\right)\right)$, where $q$ is the estimated conformal quantile from conformal prediction using validation dataset.
Given the validation dataset $\mathcal{D}_{2}$ which has $N_{2}$ data samples, conformal prediction follows the following steps: 1) Define a conformal score function $s(x,y)\in\mathbb{R}$ (smaller scores encode better agreement between x and y). 2) Compute the conformal quantile $q$ as the $\frac{\lceil(N_{2}+1)(1-\alpha)\rceil}{N_{2}}$ quantile of the validation scores $S=\{s(x_{i},y_{i})\}_{i=1}^{N_{2}}$. $\alpha\in[0,1]$ is the user-defined error rate. 3) Then for every unseen data sample $(\tilde{x},\tilde{y})$, conformal prediction forms the prediction set $\mathcal{C}(\tilde{y}|s(\tilde{x},\tilde{y})\leq q)$. Conformal prediction guarantees that the prediction set contains the true target value with a probability at least $1-\alpha$.
Gaussian Process Regression Module
@@
Existing literature of trajectory prediction models (Zhou et al. 2022, 2023; Liang et al. 2020; Zhong et al. 2022) focuses on providing point estimates of future trajectories. However, the uncertainty of the future trajectories due to the changing environment or the intrinsic intention changes of drivers can lead to significant distribution shift and overconfident trajectory prediction. Such distribution shift and overconfident prediction can greatly impact the subsequent decision-making processes, such as robust path planning (Hu et al. 2023; Kedia, Zhou, and Karumanchi 2023). Hence, we propose to consider output uncertainties from both the modeling limitation of the model $f_{\theta}$ upon current trajectory inputs and the noise inherent in the trajectory data.
@@
1 Input: a base model $f(\cdot)$ with an initialized parameter $\theta$ and a covariance function $k(\cdot)$ with an initialized parameter $\omega$, error rate $\alpha\in[0,1]$, score function $s(\cdot)$, total number of training epochs $epo$.
2 Data: training dataset $D_{1}=\{x_{i},y_{i}\}_{i=1}^{N_{1}}$, validation dataset $D_{2}=\{x^{t},y^{t}\}_{t=1}^{T_{2}}$.
@@
3 Output: well-trained base model $f_{\theta}(\cdot)$, updated statistic $\hat{q}^{1}$, score set $S={\{s\left(x^{t},y^{t}\right)\}}_{t=L-1}^{T_{2}}$ and error set $E={\{e^{t}\}}_{t=L-1}^{T_{2}}$ estimated from the validation data.